None:

Write an expression that is equivalent to three over five (3y + 15). three over fivey + 9 three over five y + 15 nine over fivey + 9 nine over fivey + 15

None:

pls help

None:

pls

Hoodmemes:

i was gonna sit here and take a question answering break srry

None:

:(

Hoodmemes:

ight

Hoodmemes:

Fist u would wanna distribute \[\frac{ 3}{ 5}\times3\rightarrow \frac{ 3 }{ 3 }\rightarrow \frac{ 9 }{ 5 }\]

None:

ok

None:

ok so we know its eather c or d

Hoodmemes:

yes

Hoodmemes:

so what answer do u think it is?

None:

d

Hoodmemes:

correct

None:

ok yay

None:

bruh that was wrong

Hoodmemes:

bto its was 9 over five plus 5

Hoodmemes:

i mean 3 over 5 plus 15 im so freaking srry